The Output Gap (also known as recessionary gap or GDP gap) is equal to $700 billion dollars. If the savings rate is 5% (MPS = 0.05) how large of a tax cut would the government have to do to raise national income enough to close the output gap?
Answer #1

Ans. To fill the output gap, increase in output required, Y = $700 billion

MPS = 0.05

=> MPC = 1-MPS = 0.95

Tax multiplier, M = -MPC/(1-MPC) = -0.95/(1-0.95) = -19

Y = M*Change in Tax

=> Change in tax = 700 billion / (-19 )

=> Change in tax = -36.8421 billion

Thus, to fill the output gap, a tax cut of $36.8421 billion in required.

* Please don’t forget to hit the thumbs up button, if you find the answer helpful.

Thank You